[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 26, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jul 26 19:55:35 UTC 2017


HSU Shift Log for Wed Jul 26 2017 15:55:28 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: Logged off at 15z per TSB instruction for NetApp, came up  
normally. On an unrelated note, JTWC is still having problem downloading  
the best track/forecasts of storms in the ATCF. Monica investigating. SHIPS  
guidance at 18z is blank for Irwin, and was blank at 12z for Greg. Mark D  
says Opah is having big problems. He printed the 18z SHIPS for Irwin off  
WCOS

N-AWIPS issues: Can not print from Tilefish. Brian M took care of it for  
us. Goes 16 data was out from 1600-1642Z in NAWIPS, similar times in  
AWIPS2. it came back right afterward but lots of black tiles in the images.

AWIPS-2/NCP issues:

Other Techincal issues: Another outage with the NRL data overnight--  
nothing since 23Z on their webpage. Data came back by noon. This is a  
recurring problem.

Miscellaneous issues: If TSB could make it a higher priority to get the G16  
SWIR loop working, that would be very beneficial to the HSU. Just the 3.9  
channel isn't as nice as Jiing's old difference SWIR, plus the color curve  
has lots of black dots still (problem fixed on vis). Thanks!
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170726/ece8845f/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list